OpenStudy (anonymous):

Medal to whoever helps me find an acceptable, satisfying answer. Kelly tells you that when variables are in the denominator, the equation one over two plus three over x equals three over four becomes unsolvable. Kelly explains, "There is a value for x that makes the denominator zero, and you can't divide by zero." Demonstrate to Kelly how the equation is still solvable and explain your reasoning.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }+\frac{ 3 }{ x }=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Could there be another value for x that allows the left side of the equal sign to be equivalent to the right side?

OpenStudy (dariusx):